OpenStudy (anonymous):

Ok so I have to solve this equation \[\sqrt{5x-3}=\sqrt{2x+3}\] Ive got it down to 5x=2x+6 how do i get 2x on the other side I know I divide.

OpenStudy (anonymous):

@actionsports

OpenStudy (campbell_st):

subtract 2x from both sides of the equation... so on the left hand side... what is 5x - 2x = ?

OpenStudy (anonymous):

its B a zero exists when 9x - 3=0 or x = 1/3 which means the curve starts at x= 1/3 the comparison graph 3square root x has a zero at the origin

OpenStudy (anonymous):

5x-2x is 3x I just asked my sister and she got me to x=2.4

OpenStudy (anonymous):

I put x into the equation and they don't add up

OpenStudy (campbell_st):

so then the equation becomes 3x = 6 solve for x.

OpenStudy (anonymous):

3x=6 /:3 x=2

OpenStudy (anonymous):

Ok thank you both!